Two adjacent angle of parallelogram (3x-51)and (5x-161)find the measure of each angle of parallelogram
Found 2 solutions by stanbon, lynnlo:
Answer by stanbon(75887)

Two adjacent angle of parallelogram (3x-51)and (5x-161)find the measure of each angle of parallelogram
----
Note: If they are adjacent they must be supplementary.
-----
Equation:
3x-51 + 5x-161 = 180
8x - 212 = 180
8x = 392
x = 49
----
3x-51 = 96 degrees
5x-161 = 84
=====================
Cheers,
Stan H.
===================

Answer by lynnlo(4176)

(3x-51)&(5x-161)
3x-51+5x-161=180
8x-212=180
8x=392
x=49
3x-51=96
5x-161=84